# Reason3D: Searching and Reasoning 3D Segmentation via Large Language Model

### v1

@@ Abstract
Recent advancements in multimodal large language models (LLMs) have shown their potential in various domains, especially concept reasoning. Despite these developments, applications in understanding 3D environments remain limited. They primarily offer textual or numerical outputs without the capability to generate dense, informative segmentation masks. This paper introduces Reason3D, a novel LLM designed for comprehensive 3D understanding. Reason3D takes point cloud data and text prompts as input to produce textual responses and segmentation masks, facilitating advanced tasks like 3D reasoning segmentation, hierarchical searching, express referring, and question answering with detailed mask outputs. Specifically, we propose a hierarchical mask decoder to locate small objects within expansive scenes. This decoder initially generates a coarse location estimate covering the object’s general area. This foundational estimation facilitates a detailed, coarse-to-fine segmentation strategy that significantly enhances the precision of object identification and segmentation. Experiments validate that Reason3D achieves remarkable results on large-scale ScanNet and Matterport3D datasets for 3D express referring, 3D question answering, and 3D reasoning segmentation tasks. Code and models are available at: https://github.com/KuanchihHuang/Reason3D.
*Corresponding author
Figure 1: We propose Reason3D, a novel LLM-based dense 3D point cloud searching and reasoning framework that can output dense segmentation masks based on textural information. Our Reason3D can handle tasks involving 1) 3D Reasoning, 2) 3D Hierarchical Searching, 3) 3D express referring, and 4) 3D QA with responding dense segmentation masks.
@@ 1 Introduction
Recently, large language models (LLMs) [41, 42, 20] have significantly enhanced their capabilities in sophisticated reasoning within the realm of natural language processing. Building on these developments, a new class of models, known as Multimodal Large Language Models (MLLMs) [14, 26, 3, 28, 54, 36, 50, 52] have emerged. These models are designed to process multiple modalities, including 2D images [28, 3], thereby enriching LLMs’ ability to interpret and understand visual inputs. However, while MLLMs demonstrate remarkable proficiency with 2D imagery, extending these capabilities to complex 3D environments remains a formidable challenge, which is critical in various technological domains, including robot navigation and embodied AI agents.
@@
Numerous studies have made significant strides in using point clouds as input tokens for 3D large language models (LLMs). Some researches [49, 33, 32, 40] primarily focused on 3D object-level understanding. Prior study 3D-LLM [17] aggregates multi-view features to enrich 3D feature comprehension and employs a subsequent LLM for 3D reasoning. Meanwhile, LL3DA [8] directly encodes 3D point clouds for scene representation and facilitates human interaction to enhance understanding.
@@
These methods integrate large language models (LLMs) with point cloud inputs to enhance 3D reasoning capabilities; however, they also encounter specific limitations. Firstly, the outputs from these models are confined to textual or numerical forms, which are inadequate for predicting dense data types, such as segmentation masks, essential for detailed spatial analysis. Secondly, these models struggle to locate or search for objects in 3D scenes based on complex or abstract concepts rather than relying on spatial relationships within the scenes.
@@
To enable 3D-based LLM models to produce segmentation masks, we can direct the LLM to generate a [SEG] token. The embedding of this token is then utilized to guide the decoder in learning to create 3D segmentation masks, similar to approaches used in 2D reasoning models [24]. However, unlike structured image data, this straightforward adaptation may encounter difficulties due to point clouds’ inherent sparsity and unstructured nature. These challenges are particularly pronounced when attempting to output masks for small objects within dense, large-scale 3D scenes.
@@
In this paper, we introduce Reason3D, a framework designed to enable reasoning and searching within 3D scenes using large language models with point cloud inputs. Unlike other methods that generate only textual and numerical outputs, Reason3D also produces 3D segmentation masks from textual inputs. To efficiently manage large-scale point clouds, we utilize a superpoint pooling layer that aggregates data into precomputed superpoints, significantly reducing complexity. Furthermore, we develop a hierarchical mask decoder to address the challenges of sparsity in extensive point clouds, such as locating a ball in a large house. This decoder employs a coarse-to-fine strategy, initially identifying a likely object-containing coarse region by instructing the LLM model to output a [LOC] embedding, which guides the learning of the region mask. This region then serves as a prior to generate a precise object mask, facilitating effective localization in complex 3D environments.
@@
Figure 1 illustrates Reason3D’s capability to handle diverse scenarios requiring advanced reasoning, searching, and question answering based on conceptual and spatial information. To validate the effectiveness of our approach, we have collected a dataset for 3D reasoning segmentation, comprising over one thousand point-instruction pairs. To ensure this dataset closely mirrors real-world applications, we annotated point clouds from Matterport3D [6] and ScanNetv2 [13] with implicit text queries that demand complex reasoning. The main contributions of this work are:
@@
  • We introduce Reason3D, a comprehensive framework for reasoning and searching within 3D scenes using extensive language prompts. This framework not only processes 3D point clouds and language prompts to generate textual outputs but also creates detailed 3D segmentation masks. Reason3D supports a wide range of tasks, including 3D expressive referring segmentation, 3D reasoning segmentation, 3D hierarchical searching, and 3D question answering.
@@
  • We establish the novel task of 3D reasoning segmentation, which involves interpreting implicit human instructions within 3D scenes, and build a dataset to evaluate this task. The capability of 3D reasoning is essential for developing truly advanced 3D perception systems.
@@
  • We develop a hierarchical mask decoder to address the sparsity of 3D point clouds effectively. This innovative approach first identifies a coarse region likely containing the object and uses this region’s probability as a prior to guide the refinement of the final mask prediction.
  • We develop a hierarchical mask decoder to address the sparsity of 3D point clouds effectively. This innovative approach first identifies a coarse region likely containing the object. Then, the decoder uses this region’s probability as a prior to guide the refinement of the final mask prediction.
@@ 2 Related Work
3D Point Cloud Segmentation. Recent advancements in point cloud segmentation [31, 23, 38, 48, 37, 44] have driven the evolution of various class-aware prediction techniques, predominantly employing UNet-like models that process data as either 3D points or voxels. Point-based methods [9, 55] refine feature extraction through aggregation mechanisms or transformer blocks, while voxel-based methods [18, 11] transform irregular point clouds into regular voxel grids for processing with dense or sparse 3D convolutional networks, enhancing segmentation accuracy and efficiency. With 3D segmentation tasks having achieved a significant level of maturity, it has become essential to develop more advanced methods for interacting with 3D segmentation systems.
@@
The 3D express referring segmentation task [7, 1] enhances interaction through human language by segmenting a 3D object according to a specific textual description. TGNN [19] employs a two-stage approach that integrates instance and textual features, computing a matching score for each instance to identify the target described in the text. Building on this, X-RefSeg3D [34] leverages vision-language cues by combining entity-related linguistic information with visual features to construct a cross-modal scene graph, which facilitates interactions based on textual and spatial relations. Similarly, 3D-STMN [45] introduces an efficient end-to-end framework that utilizes superpoints and aligns them with textual modalities, thereby enhancing their role in multimodal representation. However, while these studies make significant strides in object identification using spatial relation cues, they do not fully explore deeper reasoning capabilities. In this work, we introduce Reason3D, a novel approach that extends beyond traditional identification to incorporate advanced reasoning with 3D segmentation models, addressing complex interactions not yet tackled by existing methodologies.
@@
Large Language Model. Recent advancements in large language models (LLMs) [41, 57, 42, 20, 21] have showcased their broad generalization across diverse language tasks, thanks to training on extensive textual datasets. Through self-supervised learning techniques such as token prediction and masked token reconstruction, as well as further refinements via instruction tuning and specialized datasets, researchers have significantly enhanced the adaptability of these models to new tasks. Building on this, the remarkable reasoning capabilities of LLMs are increasingly applied in multimodal contexts. Modern models incorporate advanced architectures that integrate visual data [26, 2, 3, 28, 58, 24, 14, 47, 50, 53, 46], utilizing mechanisms such as cross-attention and image-text feature alignment to enable comprehensive multimodal understanding. This has paved the way for models that engage in visual question answering and perform complex reasoning tasks. Notably, LISA [24] introduces a specialized segmentation token into its vocabulary, decoded to generate a segmentation mask, enabling more precise and detailed reasoning capabilities.
@@
Recent efforts have extended large language models (LLMs) to include 3D data, thereby improving point cloud understanding. Point-LLM [49] interprets object-level points using LLMs, while 3D-LLM [17] enhances understanding by integrating multi-view image features with LLMs. LL3DA [8] combines textual instructions with visual interactions to improve feature extraction for more effective instruction-following. Unlike existing methods, which are limited to bounding box-level grounding, textual responses, or lack contextual reasoning, our approach enables fine-grained segmentation of precisely searched objects within 3D data.
@@
Figure 2: Examples of the annotated samples. Left is one data from Matterport3D dataset with the answer pool table. Right is one data from the ScannetV2 dataset with the answer fireplace.
@@
Dataset Collection. Given the absence of a standardized dataset for evaluating 3D reasoning segmentation, we have collected the 3D scans from indoor datasets, Matterport3D [6] and ScanNetv2 [13] and annotated them with complex text instructions and detailed 3D segmentation masks. The dataset consists of 1339 samples for training and 1145 samples for validation. Two sample data are shown in Figure 2. More details can be found in the Appendix.
@@ 4 Reason3D
Figure 3: Overview of our Reason3D framework. Initially, we utilize a point encoder to extract dense features $\mathbf{F}_{p}$ from the input scene, simplified by a superpoint pooling layer to reduce complexity. An interactor merges superpoint features $\mathbf{F}_{s}$ with a learnable query, input into a frozen LLM along with instructions to generate an output containing critical tokens, [LOC] and [SEG]. A hierarchical decoder then uses the [LOC] embedding to estimate a coarse location that likely covers the object. Finally, this estimated location integrates with the [SEG] embedding, enabling the prediction of the final segmentation masks.
We introduce Reason3D, a novel LLM-based framework for searching and reasoning within 3D point clouds, as illustrated in Figure 3. Given a 3D point cloud and a textual query describing an object of interest, our method leverages an LLM model to align point features and predict dense object segmentation masks. Section 4.1 discusses the alignment of point clouds with LLMs in the feature space. Section 4.2 introduces the proposed hierarchical mask decoder that employs a coarse-to-fine approach for generating dense segmentation masks. Section 4.3 then details the training loss of our Reason3D framework.
@@ 4.1 Alignment between LLMs and Point Cloud
Given a point cloud $\mathbf{P}\in\mathbb{R}^{N\times 6}$ consisting of $N$ points, each characterized by three colors channels (r, g, b) and three coordinates (x, y, z), we aim to extract point features and align them with existing decoder-only LLM models to facilitate scene understanding for answering complex questions.
Scene Encoder. We employ a voxelization operation on the point cloud and utilize a U-Net style backbone [16] to extract point-wise features $\mathbf{F}_{p}\in\mathbb{R}^{N\times C}$, where $C$ denotes the channel dimension. To further reduce complexity, we feed these features into a superpoint pooling layer that utilizes pre-computed superpoints [25]. This layer aggregates superpoint features $\mathbf{F}_{s}\in\mathbb{R}^{M\times C}$ through average pooling of the point-wise features within each superpoint, effectively reducing the number of points from $N$ to $M$, where $M$ represents the number of superpoints.
This reduction is crucial for managing large-scale scenes without the need to segment the point cloud into smaller segments. For instance, a single Matterport3D scene [6] contains approximately one million points—a significant challenge for existing algorithms [31] that typically require data segmentation. Our approach significantly reduces complexity and enhances processing efficiency by handling extensive data in a single pass.
Alignment with LLM. To align the superpoint features $\mathbf{F}_{s}$ with existing decoder-only LLM models, we employ an interactor $\mathcal{F}$ following Q-Former [26] to facilitate dynamic interaction between the point cloud features $\mathbf{F}_{s}$ and the learnable query $\mathbf{Q}$, resulting in an updated query $\mathbf{Q}^{\prime}=\mathcal{F}(\mathbf{Q},\mathbf{F}_{s})$. Subsequently, the updated query $\mathbf{Q}^{\prime}$ and textual instructions $\mathbf{X}_{\rm txt}$ are fed into a frozen decoder-only language model (LLM) to generate targeted responses:
@@ 4.2 Hierarchical Mask Decoder
Existing 3D scene-level LLM-based methods [17, 8] primarily produce textual or numerical outputs and cannot provide predictions for dense 3D masks. To overcome these limitations, we introduce a Hierarchical Mask Decoder that seamlessly integrates with the outputs of LLMs. This advanced integration allows our system to interpret textual instructions and generate corresponding detailed segmentation masks directly.
@@
The Hierarchical Mask Decoder predicts the masks $\mathbf{M}$, $\mathbf{M}$, using the superpoint features $\mathbf{F}_{s}$ of the input scene and prompts $\langle\mathbf{P}_{loc},\mathbf{P}_{seg}\rangle$ generated from the LLM output based on the instruction $\mathbf{X}_{txt}$:
$\mathbf{M}={\rm Decoder}(\mathbf{Q}_{h};\langle\mathbf{P}_{loc},\mathbf{P}_{seg}\rangle|\langle\mathbf{F}_{s},\mathbf{X}_{txt}\rangle),$ (2)
@@
where $\mathbf{Q}_{h}$ is a learnable object query and $\mathbf{P}_{loc}$ and $\mathbf{P}_{seg}$ denote the location and segmentation prompts, respectively. Inspired by LISA [24], which directs an LLM to output specific embedding tokens, we adopt a similar strategy. Our LLM is prompted to produce a [SEG] token. The last-layer embedding $\mathbf{h}_{seg}$ associated with the [SEG] token is transformed through an MLP projection layer $\mathcal{G}$, resulting in the segmentation prompt $\mathbf{P}_{seg}=\mathcal{G}({\mathbf{h}}_{seg})$, which is derived from the complex textual instructions to guide the final mask prediction.
@@
Moreover, for effective 3D mask prediction, especially when targeting small objects within large scenes. We propose a coarse-to-fine approach. We further introduce a special token, [LOC], to learn the coarse location that may encompass the object mask. Similar to the [SEG] token process, we refine the embedding $\mathbf{h}_{loc}$ of the [LOC] token using an MLP layer $\mathcal{G}$, resulting in $\mathbf{P}_{loc}=\mathcal{G}({\mathbf{h}}_{loc})$. To this end, we can use two prompts to guide the final mask prediction. In practice, we initially exploit a mask decoder, $\mathcal{F}_{loc}$, constructed with a transformer decoder architecture [37, 38]. Tguidingecoder uses the location prompt as the query to generate a location mask $\mathbf{M}_{loc}$:
$\mathbf{M}_{loc}=\mathcal{F}_{loc}({\mathbf{P}}_{loc},\mathbf{F_{s}}).$ (3)
Subsequently, the location mask $\mathbf{M}_{loc}$ serves as a prior to guide the final mask generation through another MLP layer $\mathcal{H}_{loc}$:
$\mathbf{M}=\mathcal{F}_{seg}({\mathbf{P}}_{seg},\mathbf{F_{s}}+\mathcal{H}_{loc}(\mathbf{M}_{loc})),$ (4)
@@
where $\mathbf{M}$ is the final mask, and $\mathcal{F}_{seg}$ shares the same architectural framework as $\mathcal{F}_{loc}$.
4.3 Training 3D LLM

### v2

@@ Abstract
Recent advancements in multimodal large language models (LLMs) have demonstrated significant potential across various domains, particularly in concept reasoning. However, their applications in understanding 3D environments remain limited, primarily offering textual or numerical outputs without generating dense, informative segmentation masks. This paper introduces Reason3D, a novel LLM designed for comprehensive 3D understanding. Reason3D processes point cloud data and text prompts to produce textual responses and segmentation masks, enabling advanced tasks such as 3D reasoning segmentation, hierarchical searching, express referring, and question answering with detailed mask outputs. We propose a hierarchical mask decoder that employs a coarse-to-fine approach to segment objects within expansive scenes. It begins with a coarse location estimation, followed by object mask estimation, using two unique tokens predicted by LLMs based on the textual query. Experimental results on large-scale ScanNet and Matterport3D datasets validate the effectiveness of our Reason3D across various tasks.
Figure 1: Overview. We propose Reason3D, a novel LLM-based 3D point cloud searching and reasoning framework that can output dense segmentation masks based on textural descriptions. Our Reason3D can handle four tasks involving 1) 3D Reasoning, 2) 3D Hierarchical Searching, 3) 3D Express Referring, and 4) 3D QA with responding dense segmentation masks.
Project lead, Corresponding author
@@ 1 Introduction
Recently, large language models (LLMs) [48, 49, 22] have significantly advanced their capabilities in sophisticated reasoning within natural language processing. Building on these advancements, a new class of models known as Multimodal Large Language Models (MLLMs) [14, 29, 3, 31, 61, 41, 57, 59, 36, 31, 3, 42] has emerged, thereby enhancing LLMs’ ability to interpret and understand visual inputs.
@@
To advance the capabilities of Multimodal Large Language Models (MLLMs) in complex 3D environments, several studies have made significant strides by using point clouds as input tokens. Some research efforts [56, 38, 37, 47] have primarily focused on 3D object-level understanding. In addition, 3D-LLM [19] aggregates multi-view features to enrich 3D feature comprehension and employs an LLM for subsequent 3D reasoning. LL3DA [10] directly encodes 3D point clouds for scene representation, facilitating human interaction to enhance understanding.
@@
These methods integrate large language models (LLMs) with point cloud inputs to enhance 3D reasoning capabilities; however, they face certain limitations. First, their outputs are restricted to textual or numerical forms, which are insufficient for predicting dense data types such as segmentation masks. Second, these models struggle to locate or identify objects in 3D scenes based on complex or abstract concepts, as they primarily rely on spatial relationships within the scenes.
@@
To enable 3D-based LLM models to produce segmentation masks, we can direct the LLM to generate a [SEG] token. The embedding of this token is then utilized to guide the decoder in learning to predict 3D segmentation masks, similar to approaches used in 2D reasoning models [26]. However, unlike structured image data, this straightforward adaptation may encounter difficulties due to the inherent sparsity and unstructured nature of point clouds. These challenges are particularly pronounced when segmenting small objects within large-scale 3D scenes.
@@
This paper introduces Reason3D, a framework that enables reasoning and searching within 3D scenes using large language models with only point cloud inputs. Unlike other methods that generate only textual and numerical outputs, Reason3D also produces 3D segmentation masks from textual queries. We first group point features into superpoints to reduce complexity and then utilize a transformer to align point features with textual instructions. These aligned features, along with query tokens, serve as the input for a pre-trained LLM. To address the challenges of segmenting small objects within extensive point clouds, such as searching for a ball in a large house, we develop a hierarchical mask decoder that employs a coarse-to-fine strategy. This strategy begins by instructing the LLM to output the [LOC] and [SEG] token embeddings. The [LOC] token guides the learning of a region mask to identify likely object-containing areas. This region mask then serves as a prior, along with the segmentation token [SEG], for generating a precise object mask, facilitating effective localization in complex 3D environments.
@@
Figure 1 illustrates Reason3D’s capability to handle diverse tasks, such as reasoning, searching, referring, and question answering. To validate the effectiveness of our approach, we collect a dataset for 3D reasoning segmentation, comprising over one thousand point-instruction pairs. These pairs are annotated from Matterport3D [6] and ScanNetv2 [13] with implicit text queries that demand complex reasoning knowledge.
The main contributions of this work are:
@@
  • We introduce Reason3D, a comprehensive framework for reasoning and searching within 3D scenes using extensive language prompts. Reason3D processes 3D point clouds and language inputs to generate both textual outputs and detailed 3D segmentation masks. It supports a wide range of tasks, including expressive 3D referring segmentation, 3D reasoning segmentation, hierarchical 3D searching, and 3D question answering.
@@
  • We establish the novel task of 3D reasoning segmentation, which involves interpreting implicit human instructions within 3D scenes, and we have built a dataset to evaluate this task.
@@
  • We develop a hierarchical mask decoder to effectively address the challenges posed by the sparsity and extensive scale of 3D point clouds. This approach first identifies a coarse region likely containing the object and uses this region’s probability as a prior to guide the refinement of the final mask prediction.
@@ 2 Related Work
3D Point Cloud Segmentation. Recent advancements in point cloud segmentation [25, 45, 55, 44, 51] have led to improved class-aware prediction techniques, predominantly employing UNet-like models that process data as either 3D points or voxels. Point-based methods [8, 62] enhance features with aggregation mechanisms or transformers, while voxel-based methods [20, 11] transform irregular point clouds into regular voxel grids for processing with 3D convolutional networks. Another line of work [35, 46, 18, 33] focuses on understanding 3D scenes with open-vocabulary inquiry. As 3D segmentation tasks mature, developing more advanced interactions with these systems using complex instruction has become essential.
@@
The 3D express referring segmentation task [7, 1] enhances interaction through human language by segmenting 3D objects based on specific textual descriptions. TGNN [21] uses a two-stage approach to integrate instance and textual features, computing a matching score to identify targets. X-RefSeg3D [39] combines linguistic and visual features to create a cross-modal scene graph for interactions based on textual and spatial relations. Similarly, 3D-STMN [52] aligns superpoints with textual inputs to enhance multimodal representation. However, while these studies make significant strides in object identification using spatial relation cues, they do not fully explore deeper reasoning capabilities. In this work, we introduce Reason3D, a novel approach that extends beyond traditional identification to incorporate advanced reasoning with 3D segmentation models, addressing complex interactions not yet tackled by existing methodologies.
@@
Large Language Model. Recent advancements in large language models (LLMs) [48, 64, 49, 22, 23] have showcased their broad generalization across diverse language tasks, thanks to training on extensive textual datasets. Through self-supervised learning techniques such as token prediction and masked token reconstruction, as well as further refinements via instruction tuning and specialized datasets, researchers have significantly enhanced the adaptability of these models to new tasks. Building on this, the remarkable reasoning capabilities of LLMs are increasingly applied in multimodal contexts. Modern models incorporate advanced architectures that integrate visual data [29, 2, 3, 31, 65, 26, 14, 54, 57, 60, 53, 36], utilizing mechanisms such as cross-attention and image-text feature alignment to enable comprehensive multimodal understanding. This has paved the way for models that engage in visual question answering and perform complex reasoning tasks. Notably, LISA [26] introduces a specialized segmentation token into its vocabulary, decoded to generate a segmentation mask, enabling more precise reasoning capabilities.
@@
Recent efforts have extended large language models (LLMs) to include 3D data for understanding point clouds. Point-LLM [56] interprets object-level points using LLMs, while 3D-LLM [19] enhances understanding by integrating multi-view image features with LLMs. LL3DA [10] combines textual instructions with visual interactions to improve feature extraction for more effective instruction-following. Unlike existing methods, which are limited to bounding box-level grounding, textual responses, or lack contextual reasoning, our approach enables fine-grained segmentation of precisely searched objects within 3D data.
@@
Figure 2: Annotated Sample Examples. (a) shows a sample from the Matterport3D dataset with the answer pool table. (b) presents a sample from the ScannetV2 dataset with the answer fireplace.
@@
Dataset Collection. Given the absence of a standardized dataset for evaluating 3D reasoning segmentation, we have collected the 3D scans from indoor datasets, Matterport3D [6] and ScanNetv2 [13] and annotated them with complex text instructions and detailed 3D segmentation masks. The dataset consists of 1339 samples for training and 1145 samples for validation. Two sample data are shown in Figure 2. More details can be found in the supplementary materials.
3D Hierarchical Searching. Building on the foundation of 3D reasoning segmentation, we can extend the task to include searching for an object within a specified location in a large-scale 3D scene based on an abstract query. For example, instead of merely finding an object to sit on, we can specify that the object should be inside a bedroom, thus precisely limiting the object’s location.
Figure 3: Overview of our Reason3D framework. Initially, we utilize a point encoder to extract point features from the input scene, which are simplified by a superpoint pooling layer to reduce complexity. An interactor merges these superpoint features with a learnable query, input into a frozen LLM along with instructions to generate an output containing specifical tokens, [LOC] and [SEG]. A hierarchical mask decoder then utilizes the [LOC] embedding to estimate a coarse location that likely covers the target object. Finally, this estimated location prior is integrated with the [SEG] embedding to enable the prediction of the final segmentation masks.
@@ 4 Reason3D
We introduce Reason3D, a novel LLM-based framework for searching and reasoning within 3D point clouds, as illustrated in Figure 3. Given a 3D point cloud and a textual query describing an object of interest, our method leverages an LLM model to align point features and predict dense object segmentation masks. Section 4.1 discusses the alignment of point clouds with LLMs in the feature space. Section 4.2 introduces the proposed hierarchical mask decoder, which employs a coarse-to-fine approach for generating dense segmentation masks. Finally, Section 4.3 details the training loss of our Reason3D framework.
@@ 4.1 Alignment between LLMs and Point Cloud
Given a point cloud $\mathbf{P}\in\mathbb{R}^{N\times 6}$ consisting of $N$ points, each characterized by three colors channels (r, g, b) and three coordinates (x, y, z), we aim to extract point features and align them with decoder-only LLM models to facilitate 3D scene understanding based on textual instructions.
Scene Encoder. We employ a voxelization operation on the point cloud and utilize a U-Net style backbone [17] to extract point-wise features $\mathbf{F}_{p}\in\mathbb{R}^{N\times C}$, where $C$ denotes the channel dimension. To further reduce complexity, we feed these features into a superpoint pooling layer that leverages pre-computed superpoints [27]. This layer aggregates superpoint features $\mathbf{F}_{s}\in\mathbb{R}^{M\times C}$ by performing average pooling on the point-wise features within each superpoint, effectively reducing the number of points from $N$ to $M$, where $M$ represents the number of superpoints.
This reduction is crucial for managing large-scale scenes without needing to divide the point cloud into smaller segments. For example, a single Matterport3D scene [6] contains approximately one million points, posing a significant challenge for existing algorithms [35], which typically require data segmentation. Our approach addresses this challenge by utilizing superpoints, enabling us to handle extensive data in a single pass.
Alignment with LLM. To align the superpoint features $\mathbf{F}_{s}$ with existing decoder-only LLM models, we employ an Interactor $\mathcal{F}$ following Q-Former [29] to facilitate dynamic interaction between the point cloud features $\mathbf{F}_{s}$ and the learnable query $\mathbf{Q}$, resulting in an output query $\mathbf{Q}^{\prime}=\mathcal{F}(\mathbf{Q},\mathbf{F}_{s})$. Subsequently, the output query $\mathbf{Q}^{\prime}$ and textual instructions $\mathbf{X}_{\rm txt}$ are fed into a frozen decoder-only language model (LLM) to generate targeted responses:
@@ 4.2 Hierarchical Mask Decoder
Current LLM-based methods for 3D scene understanding [19, 10] are limited to producing textual or numerical outputs and cannot thus predict dense 3D masks. To overcome these limitations, we propose a Hierarchical Mask Decoder (HMD) that utilizes a coarse-to-fine approach to predict segmentation tasks guided by the output of the LLM. Specifically, we first utilize a location prompt $\mathbf{P}_{\rm{loc}}$ to learn the coarse location of the mask that potentially covers the target object. This coarse location then serves as a prior for learning the object segmentation mask, guided by another segmentation prompt $\mathbf{P}_{\rm{seg}}$. We will provide a more detailed explanation later.
@@
The Hierarchical Mask Decoder predicts the segmentation masks $\mathbf{M}_{\rm{seg}}$ by utilizing the superpoint features $\mathbf{F}_{s}$ and two specific prompts $\langle\mathbf{P}_{\rm{loc}},\mathbf{P}_{\rm{seg}}\rangle$, which is based on the instruction $\mathbf{X}_{\rm txt}$:
$\mathbf{M}_{\rm{seg}}={\rm HMDecoder}(\mathbf{F}_{s};\langle\mathbf{P}_{\rm{loc}},\mathbf{P}_{\rm{seg}}\rangle|\langle\mathbf{F}_{s},\mathbf{X}_{\rm{txt}}\rangle).$ (2)
@@
To generate these two prompt features for guidance, we direct the LLM to output specific embedding tokens inspired by LISA [26]. Considering the segmentation prompt, the LLM is instructed to generate a [SEG] token. The last-layer embedding $\mathbf{h}_{\rm{seg}}$ associated with the [SEG] token is transformed through an MLP projection layer $\mathcal{G}$, resulting in the segmentation prompt $\mathbf{P}_{\rm{seg}}=\mathcal{G}({\mathbf{h}}_{\rm{seg}})$. This prompt encapsulates the features of the target object, derived from the textual instructions, to guide the final mask prediction.
@@
Moreover, to effectively target small objects within large scenes, we introduce a location token, [LOC], which the LLM is instructed to generate. This token learns a coarse location that may potentially encompass the object mask, serving as a prior feature to enhance the accuracy of the final segmentation results. Similar to the [SEG] token process, we refine the embedding $\mathbf{h}_{\rm{loc}}$ of the [LOC] token using an MLP layer $\mathcal{G}$, resulting in the location prompts $\mathbf{P}_{\rm{loc}}=\mathcal{G}({\mathbf{h}}_{\rm{loc}})$. To this end, we can use these two prompts to guide the final mask prediction. In practice, we first exploit a region decoder, $\mathcal{F}_{\rm{loc}}$, built on a transformer decoder architecture [44, 45]. The location prompt $\mathbf{P}_{\rm{loc}}$ serves as the query, while the superpoint features $\mathbf{F_{s}}$ act as key and value to generate the location mask $\mathbf{M}_{\rm{loc}}$:
$\mathbf{M}_{\rm{loc}}=\mathcal{F}_{\rm{loc}}({\mathbf{P}}_{\rm{loc}},\mathbf{F_{s}}),$ (3)
where the location mask indicates the probability of the region potentially covering the target object.
After that, we use an MLP layer $\mathcal{H}_{\rm{loc}}$ to encode the location mask $\mathbf{M}_{\rm{loc}}$, serving as a feature prior, which is then combined with the point features. Subsequently, we make the final mask generation based on the segmentation prompt and the integrated point features:
$\mathbf{M}_{\rm{seg}}=\mathcal{F}_{\rm{seg}}({\mathbf{P}}_{\rm{seg}},\mathbf{F_{s}}+\mathcal{H}_{\rm{loc}}(\mathbf{M}_{\rm{loc}})),$ (4)
@@
where $\mathbf{M}_{\rm{seg}}$ is the final mask, and $\mathcal{F}_{\rm{seg}}$ shares the same architectural framework as $\mathcal{F}_{\rm{loc}}$.
4.3 Training Reason3D